### v1

# Examining dark neutron decay using f-mode gravitational waves

Abstract
In this study, the impact of neutron decay into dark matter and various dark matter self-interaction strengths on neutron star properties have been explored. Using the quark-meson coupling (QMC) model for nucleon-only equations of state (EoSs), the effects of different matter compositions have been compared, including strange matter and self-interacting dark matter. The results demonstrate that increasing DM-DM self-repulsion stiffens the EoS, influencing the mass-radius relationship and stability of neutron stars.
Furthermore, fundamental mode (f-mode) oscillations have been analyzed, which serve as a diagnostic tool for probing neutron star interiors. Both first- and second-order f-mode frequencies follow universal relations, reinforcing their applicability for constraining dense matter properties. It has been shown that neutron stars composed of nucleons-only and self-interacting dark matter exhibit a universal behavior in damping time and angular frequency, whereas strange matter and non-self-interacting dark matter deviate from this trend.
These findings highlight the significance of gravitational wave detections in investigating neutron star structure, particularly with next-generation observatories such as the Einstein Telescope and Cosmic Explorer. This study provides a pathway for testing dark matter models through gravitational wave asteroseismology, bridging astrophysical observations and fundamental particle physics.
1 Introduction
The hypothesis of self-interacting dark matter (SIDM) has gained prominence due to its potential to address discrepancies within the standard collisionless cold dark matter (CCDM) paradigm, alongside its theoretical plausibility within extensions of the Standard Model. Specifically, the CCDM model faces challenges such as the observed flat density profiles in dwarf galaxy cores, contrasting with the cuspy profiles predicted by simulations [1, 2], the overabundance of simulated satellite galaxies compared to observed counts [1, 3, 4], and the ”too big to fail” problem, wherein simulations predict massive, star-forming dwarf galaxies that remain undetected [5].
While baryonic interactions and statistical fluctuations have been proposed as potential resolutions to galactic anomalies, SIDM offers a compelling alternative capable of addressing these issues, and the neutron decay hypothesis provides a specific mechanism for dark matter production. Enhanced self-scattering rates in high-density regions, resulting from SIDM, could effectively flatten dwarf galaxy cores. The required self-interaction cross-sections, within the range of 0.1cm10 cm, where is the self-interaction cross-section and the dark matter particle mass are, have been extensively studied [6, 7, 8, 9, 10].
Additionally, the neutron lifetime anomaly, where beam experiments yield longer lifetimes than bottle experiments [11, 12, 13, 14], suggests a potential decay channel into undetected dark sector particles [15, 16, 17, 18]. This has motivated the exploration of neutron decay into dark matter as a potential resolution. The neutron decay scenario, proposed by Fornal and Grinstein [17], posits that neutrons decay into a dark fermion (), potentially explaining the observed neutron lifetime discrepancy. This hypothesis implies that neutron stars, composed primarily of neutrons, could serve as natural laboratories for studying dark matter interactions. If neutrons decay into dark matter within these stars, the accumulated dark matter could significantly influence their properties.
The fundamental mode (f-mode) is a type of non-radial oscillation in compact stars, primarily governed by the star’s bulk properties—such as mass, radius, and internal composition. These oscillations generate gravitational waves, and their frequency and damping time are sensitive to the equation of state (EoS) of the stellar interior. As such, the f-mode serves as a powerful diagnostic tool for probing the internal structure and composition of neutron stars and, more generally, for constraining the nature of dense matter under extreme conditions.
There are several promising avenues for measuring f-mode [19, 20, 21, 22] characteristics in compact stars. For instance, third-generation gravitational wave detectors like the Einstein Telescope [23] and Cosmic Explorer [24] are expected to achieve the sensitivity required for detecting f-mode signals from astrophysical sources [25].
Furthermore, the damping time and frequency of the f-mode are also linked to the star’s properties, such as mass, compactness, and moment of inertia, quantities that could be inferred from spin-orbit coupling effects in binary pulsar systems [26]. These interdependencies allow the f-mode to act as a bridge between observable gravitational wave signatures and the otherwise inaccessible interior physics of compact stars.
In this study, well-established universal relations for neutron stars (NSs) involving f-mode properties—relations that have been foundational in the field due to their near-independence from the specific choice of nuclear EoS, have been revisited [27, 28]. This investigation has been extented to isolated compact stars composed of dark matter (DM), modeled as static, spherically symmetric objects with EoSs for both fermionic and bosonic dark matter particles.
In essence, f-mode oscillations provide a unique window into the microphysics of compact stars. For dark matter candidates, this approach opens a novel and testable pathway to constrain their properties via gravitational wave asteroseismology.
This paper has been divided into different sections. The layout of this study is as follows: Section 2 provides theoretical possibility of neutron decay into dark matter particles, examining potential mechanisms and astrophysical constraints. Section 3 provides a comprehensive overview of neutron stars, detailing their formation, key properties, and observed phenomena. Section 4 focuses on the crucial concept of the Equation of State (EoS) governing the ultradense matter within neutron stars, exploring theoretical models and their implications. Building upon this, Section 5 discusses the fundamental structural equations that describe the equilibrium of neutron stars under their immense gravity. Section 6 further elaborates on the Equation of State, potentially exploring specific aspects or advanced considerations. Section 7 analyzes the intriguing phenomenon of universalities observed in neutron star properties, highlighting equation-of-state independent relationships. This section will be further subdivided, with Section 7.1 specifically focusing on the damping time of neutron star oscillations and its physical mechanisms. Section 7.2 will then examine the mass-scaled angular frequency of rotating neutron stars and its significance. Finally, Section 8 will conclude by focusing on the f- mode oscillation frequencies of neutron stars and their importance for probing their internal structure and for gravitational wave astronomy.
2 Neutron decay into dark matter
There remains an ongoing discrepancy between neutron lifetime measurements obtained using beam and bottle methods. Beam experiments, which detect proton decay, consistently yield lifetimes approximately 1% longer than those measured in bottle experiments, where the decay products are not directly observed. Advances in ultracold neutron studies have refined bottle measurements, yet beam experiments continue to produce lifetimes that remain about eight seconds longer [29, 18, 30, 31, 32, 33, 34, 35, 36, 37, 36].
To address this anomaly, Fornal and Grinstein [15, 16, 17] proposed a dark decay channel wherein neutrons decay into dark sector particles,
including a dark fermion (), alongside the standard mode.
This hypothesis posits that weakly interacting dark matter particles remain undetected in beam experiments but contribute to the total decay rate in bottle experiments. This alternative to neutron-mirror neutron oscillation [15, 16] necessitates a dark fermion with a mass (mχ) nearly degenerate with the neutron
Within the range [15, 16]. The accompanying boson cannot be a photon [38, 39]. Furthermore, this dark decay hypothesis has been suggested as a potential resolution to the reactor anti-neutrino anomaly [39]. Notably, it imposes constraints on neutron star properties. If these dark fermions are non-interacting, the maximum neutron star mass is predicted to be below 0.7 M⊙ [40, 41, 42, 43, 44], contradicting observed neutron star masses exceeding 2 M⊙ [45, 46, 47, 48]. Therefore, a repulsive self-interaction among dark fermions is required to reconcile this discrepancy.
This study investigates the implications of this dark decay channel on neutron star properties. This study takes forward the previous findings and constrains the properties of the dark matter using the gravitational wave, which can be very crucial in determining the interior of the neutron star.
3 Neutron Stars
Neutron stars are extremely compact objects that exist in the universe and exhibit an extreme range of energy densities. In fact, black holes are the only objects that are more compact than neutron stars. The energy densities at the outer crust are comparatively lower, and this part plays a crucial role in determining the radius of the neutron stars At the core, neutron stars contain energy densities several times the energy density of normal nuclear matter. The composition of matter at such high energy densities has been a puzzle for physicists. There have been speculations proposed by different physicists suggesting such high energy densities at the core of neutron stars may contain nuclear matter, hyperons, or even strange matter [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82]. But it has been widely accepted that mostly 90% of it is made of neutrons. This neutron-rich environment renders neutron stars ideal astrophysical laboratories for examining the proposed neutron decay into dark matter (DM) scenarios.
The hypothesis that neutrons may decay into DM suggests that a substantial number of dark matter particles may accumulate within neutron stars, potentially leading to observable alterations in their macroscopic properties [83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 43, 94, 95, 96, 97, 98, 99, 100, 101]. To model the nuclear matter inside neutron stars with precision, the equation of state (EoS) based on the quark-meson coupling (QMC) model has been used [102, 103, 104, 105]. This model is rooted in a relativistic mean field. This framework, grounded in relativistic mean-field theory, effectively represents nucleon interactions through meson exchange. In this study, I deliberately limit the hadronic composition to nucleons alone, omitting hyperons and strange quark matter. This simplification allows us to focus specifically on the effects of neutron-to-dark matter decay within a well-defined nuclear setting, facilitating a more precise analysis of its influence on neutron star properties. Through this approach, I systematically examine how dark matter accumulation affects observable features of neutron stars, offering valuable insights into dark matter’s fundamental characteristics.
4 Equation of State
To determine the equation of state, the QMC model has been adapted, which has shown promising results in [82, 106, 107, 108]. The quark-meson coupling (QMC) model was first introduced by Guichon and later expanded upon by Guichon, Thomas, and their collaborators. In this framework, nucleons are modeled as clusters of three quarks confined within an MIT bag, emphasizing their internal structure—unlike other models that treat nucleons as point-like particles [109]. The interactions between baryons arise from meson exchange, with the mesons coupling dynamically to the confined quarks. A key feature of this model is the strong scalar mean field, which induces significant modifications in the properties of bound baryons. As proposed by Fornal and Grinstein [17, 16], if neutron stars decay into dark matter. Dark matter must be taken into account when formulating an equation of state. At ()-equilibrium, the chemical equilibrium equations governing the composition of matter are:
The energy density of the system, including the dark matter, is given by:
where is the mass of corresponding particle (subscripted). The nucleon effective mass, , given by
Here is the -nucleon coupling constant in free space calculated in the model. The pressure () is calculated using
The equation of state (EoS) describes the relationship between pressure () and energy density (). A generic form including dark neutron contributions can be written as:
where and represent the baryonic and dark neutron contributions, respectively. A common parameterization for dark matter influence is given by:
where is the dark matter equation of state parameter. The nucleons-only EoS is taken from [105] and strange matter EoS is drive based on MIT bag model and the EoS used is given by
where the energy density of the strange matter and is the bag constant, taken to be 1014gm/fm3.
5 Structural equations
Neutron stars are dense astrophysical objects shaped by intense nuclear and gravitational forces. The potential presence of dark matter within these stars may influence their structural characteristics, resulting in changes to their equation of state (EoS). This study examines methods for integrating dark neutron components into the conventional neutron star EoS and solving the Tolman-Oppenheimer-Volkoff (TOV) [110, 111] equations to assess their impact
TOV equation
where is the pressure, is the energy ensity, is the gravitational potential, is the mass, and is the radial coordinate
Perturbation Equations for f-mode Oscillations
where represents the radial displacement function, is the perturbation of the gravitational potential, is the angular frequency of oscillations, and is the spherical harmonic index.
Boundary Conditions and Numerical Integration
To solve these equations, boundary conditions are applied:
At the center:
(5.6) -
At the surface:
The equations are integrated numerically. The eigenvalues of the system correspond to the f-mode frequencies.
6 Relationship between energy density and pressure
Figure 1 illustrates the equation of state of different models selected in this study. For the nucleon-only equation of state, the QMC model is selected. The equation of state describing neutron decay into dark matter is presented for different DM-DM self-interactions. Various equations of state are considered, each corresponding to a different DM-DM self-interaction. For comparison, strange matter EoS is also given. Fig. 1 shows that the nucleon-only equation of state is the stiffest among the selected models, whereas increasing DM-DM self-repulsion makes the equation of state stiffer as the self-interaction increases the equation of state neutron decay into DM comes closer to the nucleon-only EoS. Here the self-interaction of dark matter is thought to resemble the neutron- coupling, represented by (). It is defined as:
where is the coupling constant, and is the mass of the interchange particle. Various values of are selected in this study, where the number next to represents the strength of the DM-DM particle interaction. A higher value indicates stronger DM-DM self-interaction, influencing the overall dynamics and structural properties of the system.
The strange matter EoS poorly represents the properties of nuclear matter at lower energy densities. Therefore, below 500 MeV/fm3, the strange matter EoS matches with the nucleon-only EoS [105, 107].
Fig. 2 represents the mass and radius of neutron stars based on different EoSs. Ref. [115, 116, 117] report the observation of neutron stars of a mass of exceeding 2 M⊙, indicating that any viable model must support a maximum neutron star mass above this threshold 2 M⊙. The nucleon-only EoS produces a neutron star of maximum mass greater than 2 M⊙, the strange matter EoS produces a neutron star of maximum mass close to 2 M⊙, while different EoSs based on different DM-DM self-interactions produce a neutron star of less than 2 M⊙. The non-self-interactive ( = 0 fm2) DM EoS produces a neutron star of mass close to 0.7 M⊙ which is well below the acceptable value. When the DM-DM self-interaction . The EoS produces a neutron star of 2 M⊙. Therefore, acceptable self-interaction strength is 26 fm2.
7 Universalities
The universalities refer to empirical relations that link the fundamental oscillation modes of compact objects, such as neutron stars and dark stars, to their macroscopic properties like mass, radius, and compactness. These universal relations are valuable because they provide a model-independent way to infer stellar characteristics from gravitational wave observations [118, 113]. Traditional neutron stars follow well-established scaling laws for f-mode frequencies and damping times, but recent studies show that dark matter admixed compact stars deviate from these relations, requiring new frameworks to describe their oscillations [119, 120]. By leveraging f-mode gravitational wave signals, astrophysicists aim to refine equations of state and probe exotic forms of matter beyond the Standard Model.
7.1 Damping time
Reference [27] introduced a first-order universal relation describing the damping time () of f-mode oscillations. Their estimate was based on the quadrupole radiation formula, and the corresponding result is illustrated in Figure 3.
This equation expresses the characteristic damping time () of an oscillating astrophysical object, showing its relationship to the oscillation energy and the power emitted through gravitational waves (GW). The notation suggests that the damping time is proportional to the total oscillation energy divided by the power emitted as gravitational waves. The duration of oscillations in an astrophysical system depends on how efficiently it emits gravitational waves, with larger stars generally retaining their oscillations longer. Building on this, Lioutas and Stergioulas [121] later refined the relation by including higher-order corrections. The results illustrated in Fig. (3) demonstrate the first and second-order damping times for neutron stars, confirming that all selected equations of state (EoSs)—including nucleons-only matter, strange matter, and neutron decay into dark matter—exhibit a universal behavior. This universality is evident from the precise overlap of their plots for both first- and second-order equations. Furthermore, Fig. (3) highlights that dark matter, when incorporated into the neutron star’s composition, consistently exhibits universal behavior. This universality holds true regardless of the strength of dark matter’s self-interaction, as indicated by recent findings [44]. These results provide critical insights into neutron star dynamics across various theoretical models, offering potential constraints on the properties of dark matter and its interaction mechanisms within compact astrophysical objects. Additionally, the second-order damping time exhibits a much steeper decline compared to the first-order, with both damping times decreasing as the compactness of the neutron star increases.
The relationship between normalized damping time and neutron star mass, as depicted in Fig. (4), highlights that heavier neutron stars experience reduced damping time. Interestingly, while nucleons-only EoS and all self-interacting dark matter EoSs follow a universal trend and nearly overlap, strange matter EoS and neutron decay into dark matter EoS without self-interaction diverge from this universality.
7.2 Mass-scaled angular frequency
Figures (5) and (6) which are calculated using Ref. [122], illustrate the relationship between mass-scaled angular frequency and both compactness and effective compactness. These figures provide valuable insights into how the angular frequency varies with neutron star structure, highlighting key trends influenced by mass and gravitational binding. As shown in Fig. (5), the angular frequency of neutron stars increases with their compactness, suggesting that more compact neutron stars may emit stronger gravitational waves, potentially detectable. However, the strange matter EoS and dark matter EoS without self-interaction appear to deviate from the trend observed in nucleons-only and self-interacting dark matter EoSs. Notably, nucleons-only and self-interacting dark matter EoSs exhibit a universal behavior, aligning closely across different models.
Figure (6) illustrates the universal behavior observed across the selected equations of state (EoSs), including strange matter and neutron decay into dark matter EoS. The plot demonstrates that, despite differences in composition, these EoSs follow a consistent trend, reinforcing the idea that the mass-scaled angular frequency maintains a predictable relationship with effective compactness (). This universal pattern suggests that key physical properties governing neutron star oscillations remain largely unaffected by variations in the underlying matter composition, except for minor deviations. The alignment of different models further supports the robustness of this behavior across a wide range of astrophysical scenarios.
8 F-mode frequency
The frequency of an f-mode is highly sensitive to the EoS of the ultra-dense matter within the neutron star [112, 113, 114]. A stiffer EoS (less compressible) generally predicts higher f-mode frequencies. Conversely, a softer EoS (more compressible) predicts lower f-mode frequencies. The acceptable values of f-mode frequency for compact stars, such as neutron stars, typically range between 1 and 1–3 kHz. These frequencies are significant because they fall within the detection range of advanced gravitational wave detectors like the Einstein Telescope and Cosmic Explorer [23]. Figure (7) illustrates that the f-mode frequency of the selected models adheres to the universal relation, demonstrating consistency across different astrophysical scenarios. Notably, both frequency ranges fall within the detection capabilities of the Einstein Telescope and Cosmic Explorer, highlighting their potential for observational verification. Furthermore, for neutron decay into dark matter, variations in dark matter self-interaction strength do not disrupt the universal trend. Regardless of the DM-DM self-interaction parameter, the f-mode frequency remains well within the expected range, reinforcing the robustness of the underlying physical principles governing these oscillations. This consistency suggests that such models could be effectively probed through next-generation gravitational wave observatories.
9 Conclusion
In conclusion, this study highlights the impact of different equations of state (EoSs) on neutron star properties, including mass, radius, damping time, and oscillation frequencies. The findings confirm that nucleon-only EoSs produce the stiffest models, while increasing dark matter (DM) self-repulsion strengthens the equation of state. Notably, neutron decay into dark matter aligns more closely with nucleon-only EoSs as self-interaction increases.
The study further reinforces the existence of universal relations in neutron star oscillations. Both the first- and second-order damping time equations demonstrate consistency across selected EoSs, with strange matter and non-self-interacting dark matter exhibiting deviations. Similarly, mass-scaled angular frequency correlates with compactness, suggesting that more compact neutron stars may emit stronger gravitational waves, making them promising candidates for detection.
Additionally, the f-mode frequency results indicate that all selected models follow a universal trend, remaining within the detection range of next-generation observatories such as the Einstein Telescope and Cosmic Explorer. This consistency supports the viability of probing neutron star interiors through gravitational wave observations, emphasizing the importance of understanding the role of dark matter interactions in dense astrophysical objects.
The results of this study enhance our understanding of neutron star behavior and provide valuable insights for future gravitational wave detections, opening avenues for further investigations into exotic matter compositions and their astrophysical implications.
Acknowledgments
I sincerely thank Professor Anthony Thomas (University of Adelaide) for his invaluable corrections and insightful suggestions, which have significantly enhanced the clarity and accuracy of this work. His expertise and guidance were instrumental in refining our analysis, and I deeply appreciate his suggestion to this study.